Stiller Company had the following information for its three intangible assets.

  1. Patent: A patent was purchased for $200,000 on June 30, 2018. Stiller estimated the useful life of the patent to be 15 years. On December 31, 2020, the estimated future cash flows attributed to the patent were $170,000. The fair value of the patent was $150,000.
  2. Trademark: A trademark was purchased for $10,000 on August 31, 2019. The trademark is considered to have an indefinite life. The fair value of the trademark on December 31, 2020, is $5,000.
  3. Goodwill: Stiller recorded goodwill in January 2019, related to a purchase of another company. The carrying value of goodwill is $60,000 on December 31, 2020. On December 31, 2020, the segment for which the goodwill relates had a fair value of $1,160,000. The book value of the net assets of the segment (including goodwill) is $1,200,000.

Note: Round each of your answers to the nearest whole dollar.

a. Classify each of the intangible assets above as a finite life intangible or an indefinite life intangible.

b. Determine the carrying value of each asset on December 31, 2020, prior to testing for impairment, assuming that the company uses the straight-line method to amortize intangible assets, and no impairment was reported prior to 2020.

Carrying value of patent, Dec. 31, 2020, before impairment testing
Carrying value of trademark, Dec. 31, 2020, before impairment testing
Carrying value of goodwill, Dec. 31, 2020, before impairment testing

c. Test each asset for impairment assuming that the qualitative assessment indicated that further impairment testing was warranted. Determine the carrying value of each asset on December 31, 2020, after impairment testing.

Carrying value of patent, Dec. 31, 2020, after impairment testing
Carrying value of trademark, Dec. 31, 2020, after impairment testing
Carrying value of goodwill, Dec. 31, 2020, after impairment testing
Answer #1
Part a Part b Part c
Impairment Carrying Value after impairement
Patent Finite life 199998 200000-170000=30000 170000
Trademark Indefinite life 10000 10000-5000=5000 5000
Goodwill Indefinite life 60000 Fair Value of segment=1160000 40000
Book value of assets including goodwill=1200000
Value of goodwill = 1200000-1160000= 40000
Impairment=60000-40000=20000